### v1

@@ 5.4 Numerical Simulations of NNs: Time Evolution
Figure 5 was created using the trained StNN and FFNN to take an initial state and autoregressively advance the solution by $\Delta t$. The output at each time stamp was reinserted into the NNs to estimate the solution $k\Delta t$ to predict time-advanced states. This iterative mapping could produce a prediction for the future state as far into the future as desired. More specifically, figure 5 shows states mapping from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ to predict Lorenz solution 600-time steps into the future from a given initial state. The performance of the StNN was then compared with the FFNN to approximate the future dynamics of the system. The evolution of two randomly chosen trajectories is predicted using the StNN and FFNN as shown in figure 5. Both networks show remarkable accuracy in predicting highly chaotic and non-linear dynamics to map states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$. To elaborate on this further, we also compare with SINDy and HAVOK, showing the time evolution of the individual components within the states ${\bf x}_{k}$ against the NNs prediction in the Section 5.5.
5.5 Comparison of StNN, FFNN, SINDy, and HAVOK
In this section, we utilize the StNN associated with the lowest validation MSE, where $p=6$ to compare its performance against the FFNN, SINDy, and HAVOK models, focusing on accuracy and flop counts. For this comparison, we reference a benchmark simulation derived from the Lorenz system as outlined in [13]. For these simulations, we generated trajectories using the Lorenz equations with the same parameters discussed in Section 4.2 that were used to simulate the StNN and FFNN. The system is simulated up to $T=50$ with a time step of $dt=0.01$, resulting in a dataset containing $5000$ time steps. Initial states $(x,y,z)=(0,1,20)$ were used to simulate the system, and the generated data was split into training and testing sets, with $80\%$ allocated for training and a subset of $600$ time steps for testing. We utilized the PySINDy Python package [17] to simulate the SINDy model and the PyDMD Python package [19, 24] to simulate the HAVOK model. Next, we created a 600-step random trajectory for the test, and we provided the FFNN, StNN, SYNDy, and HAVOK with the trajectory’s first initial condition. The 600-step trajectory was then iteratively predicted by running each model 600 times. The predicted values were compared with the actual values to calculate the MSE for 600 steps across three position values. Table 4 and Figure 6 show the accumulated MSE across $600$ steps for all the models. The SINDy operator displayed interpretability by presenting an explicit mathematical representation of the dynamics in the form of sparse equations. Its minimal processing overhead makes it ideal for systems with simple, well-defined governing equations. However, the SINDy model has a higher MSE accumulation compared to StNN, especially after 100 time steps.
@@ 5.5 Comparison of StNN, FFNN, SINDy, and HAVOK
Figure 6: Comparison of predicted and true trajectories along the x, y, and z axes over 600 time steps for different models. The FFNN (a) and StNN (b) models are the trained machine learning models used for trajectory prediction, while SINDy (c) and HAVOK (d) serve as classical algorithm baselines for comparison. The blue solid lines represent the true trajectory, while the red dashed lines indicate the predicted trajectory. FFNN and StNN show strong alignment with the true trajectory, whereas SINDy exhibits deviations at later time steps. The HAVOK model fails to maintain stability, leading to extreme numerical divergence. This comparison highlights the effectiveness of FFNN and StNN in capturing system dynamics compared to traditional methods.
Table 4: Comparison of MSE, flop counts, and parameters among FFNN, StNN, SYNDy, and HAVOK
ML or Data-driven Accumulated MSE Inference flop Parameters
Algorithm at 100 steps counts
FFNN $0.288$ $3960$ $2073$
StNN $0.566$ $932$ $408$
SINDy $2.452$ $209$ $7$
HAVOK $16.127$ $544$ $1825$
In contrast, the StNN offered better performance for complex, nonlinear dynamics. While it required a less computationally intensive training phase than FFNN, the StNN achieved a smoother and lower cumulative MSE compared to the SINDy and HAVOK operators over the 100-time step trajectory as listed in Table 4.
In terms of computational complexity, SINDy is lightweight, using substantially less memory and flops than StNNs. However, the StNN was shown to be more resilient in precisely simulating the chaotic Lorenz system.
In conclusion, while the FFNN surpasses the StNN in terms of MSE, the StNN has a clear edge in terms of computational efficiency, with fewer flops and weight counts. These qualities make the StNN an attractive choice for use in resource-constrained contexts where efficiency takes importance over absolute error minimization.
6 Conclusions
@@
In this paper, we proposed a low-complexity structured neural network (StNN) for modeling and predicting the evolution of dynamical systems. Our approach uses the Hankel operator to give a structured and computationally efficient alternative to conventional neural networks and data-driven techniques like LEADs, SINDy, and HAVOK. According to numerical simulations based on the Lotka-Volterra model and the Lorenz system, the proposed StNN outperformed other methods in terms of decreasing computer complexity while retaining accurate long-term trajectory predictions. Our findings show that the structured nature of the Hankel operator-based neural network considerably decreases the number of parameters and flop counts while increasing the efficiency of the StNN when compared to conventional techniques. Furthermore, comparisons to baseline approaches, such as FFNN and conventional operator-based techniques, demonstrate StNNs’ promise for real-time applications in highly non-linear and chaotic dynamical systems.
@@
Future work will include expanding the StNN framework to higher-dimensional dynamical systems, and utilize the Hankel operator defined through the observation of the state to efficiently solve PDEs.

### v2

@@ IV-B Numerical Setup for the Chaotic Lorenz System
Figure 3 was created using the trained StNN and FFNN to take an initial state and autoregressively advance the solution by $\Delta t$. The output at each time stamp was reinserted into the NNs to estimate the solution $k\Delta t$ to predict time-advanced states. This iterative mapping could produce a prediction for the future state as far into the future as desired. More specifically, figure 3 shows states mapping from ${\bf x}_{k}$ to ${\bf x}_{k+1}$ to predict Lorenz solution 600-time steps into the future from a given initial state. The performance of the StNN was then compared with the FFNN to approximate the future dynamics of the system. The evolution of two randomly chosen trajectories is predicted using the StNN and FFNN as shown in figure 3. Both networks show remarkable accuracy in predicting highly chaotic and non-linear dynamics to map states from ${\bf x}_{k}$ to ${\bf x}_{k+1}$. To elaborate on this further, we also compare with SINDy and HAVOK, showing the time evolution of the individual components within the states ${\bf x}_{k}$ against the NNs prediction in the Section IV-B1. A short-term comparison table with StNN, FFNN, DMD, SiNDY, and HAVOK followed by a paragraph about it, should go here.
IV-B1 Comparisons and Predictions of StNN, FFNN, DMD, SINDy, and HAVOK
In this section, we utilize the StNN associated with the lowest validation loss, where $p=6$ to compare its performance against the FFNN, DMD, SINDy, and HAVOK models, focusing on accuracy and flop counts. For this comparison, we use the benchmark simulations of the Lorenz system from DMD, SINDy, and HAVOK based on [16].
TABLE III: Comparison of training error, number of learnable parameters, and training time among StNN, FFNN, DMD, SINDy, and HAVOK.
Technique Training Error (12) Parameters Training Time
StNN $1.26\times 10^{-6}$ 536536536 27 min.
FFNN $0.84\times 10^{-6}$ 207320732073 72 min.
DMD $2.48\times 10^{2}$ 151515 0.47 sec.
SINDy $1.29\times 10^{-6}$ 777 6 sec.
HAVOK $8.59\times 10^{-4}$ 182518251825 0.54 sec.
Table III highlights the training error, number of learnable parameters, and training time among StNN, FFNN, DMD, SINDy, and HAVOK of the Lorenz system. The proposed StNN reduces the number of parameters to just 536 and also the training time by more than 75%, with an error order $10^{-6}$ in 27 minutes compared to FFNN with 2073 parameters and a longer training time of 72 minutes. Compared to FFNN and HAVOK, the proposed StNN is characterized by having the minimum number of parameters. Even though DMD requires the shortest training time compared to StNN, FFNN, SINDy, and HAVOK, it results in the highest error, showing that it does not accurately reflect the chaotic dynamics of the Lorenz system. The HAVOK also trains quickly, but its higher training error compared to StNN reflects a trade-off between model simplicity and how well they fit the training data. Interestingly, SINDy achieves the lowest parameters, although it does not show prominent results for long-term behavior as of StNN shown in Table IV.
[…]
TABLE IV: Predictions for the testing accuracy, inference flops, and inference time among StNN, FFNN, DMD, SINDy, and HAVOK.
Technique Testing Inference Inference Time (ms)
Accuracy (12) flops
StNN $2.43\times 10^{-5}$ 884 1.30
FFNN $0.99\times 10^{-5}$ 396039603960 0.36
DMD $2.09\times 10^{2}$ 515151 0.19
SINDy $0.48\times 10^{-5}$ 209209209 1.66
HAVOK $1.90\times 10^{-3}$ 544544544 0.20
Based on Table IV, the proposed StNN requires fewer flops, i.e., 78% reduced flops compared to FFNN with an accuracy order $10^{-5}$. Also, the accuracy in StNN is higher than that in HAVOK and much higher than in DMD. The inference time of StNN is lower than that of SINDy, but not as low as HAVOK and DMD. On the other hand, DMD has a higher error compared to the StNN and other models due to the linear approximation of the nonlinear model. Finally, we note here that FFNN, DMD, SINDy, and HAVOK benefit from PyTorch’s highly efficient, GPU-accelerated operations. But, when we set $p=1$ in Table II, the proposed StNN achieves an inference time of 0.238 ms while standing out among FFNN, DMD, SINDy, and HAVOK. Also, based on the proposed StNN codes in StNN-Dynamical-Systems, StNN currently has not been implemented with low-level parallel optimization and has not been benefited by PyTorch’s libraries. Certainly, the StNN codes can be parallelized, which will further minimize both flops and inference time compared to optimized PyTorch libraries in FFNN, DMD, SINDy, and HAVOK. Thus, compared to FFNN, DMD, SINDy, and HAVOK, StNN is a lightweight and low-complexity neural network suitable for modeling dynamical systems, making it particularly ideal for chaotic systems with well-defined governing equations.
@@ IV-B1 Comparisons and Predictions of StNN, FFNN, DMD, SINDy, and HAVOK
Figure 4: Comparison of predicted and true trajectories along the x, y, and z axes over 500 time steps for different models. The FFNN (a) and StNN (b) models are the trained machine learning models used for trajectory prediction, while SINDy (c), HAVOK, (d) and DMD (e) serve as classical algorithm baselines for comparison with Python packages in [47, 48, 49]. The blue solid lines represent the true trajectory, while the red dashed lines indicate the predicted trajectory. The StNN followed by FFNN shows strong alignment with the true trajectory prediction, whereas DMD, SINDy, and HAVOK exhibit deviations. Among SINDy, DMD, and HAVOK, the DMD model fails to maintain stability, leading to extreme numerical divergence. This is since the DMD model is a simple linear operator by construction; it struggles to capture the nonlinear dynamics. In summary, this comparison shows that StNN is significantly more accurate in long-term predictions than conventional methods like DMD, SINDy, and HAVOK.
As shown in Fig. 4, for long-term predictions, SINDy accumulates a higher error compared to the StNN, particularly after 300 time steps. Thus, the SINDy tends to diverge from the predicted trajectory, while the StNN model continues to provide predictions that closely follow the trajectory for the given periods. Moreover, since the DMD model is a simple linear operator by construction, it struggles to capture the nonlinear dynamics of the Lorenz system. This results in a higher error of SINDy compared with StNN, as shown in Table IV and illustrated in Fig. 4. Thus, the proposed StNN shows better performance while capturing complex, chaotic, and nonlinear dynamics of the Lorenz system compared to SINDy. Furthermore, StNN required a less computationally intensive training phase than FFNN with a smoother and lower cumulative loss compared to DMD, SINDy, and HAVOK, as shown in Fig. 4. In terms of computational complexity, SINDy is lightweight, using fewer flops than StNNs. However, the StNN was shown to have lower inference time compared to SINDy and the best accuracy compared to SINDy, DMD, and HAVOK while precisely simulating the chaotic Lorenz system for long-term prediction. In conclusion, the StNN outperforms all models in testing accuracy and inference time for long-term predictions of the chaotic Lorenz system. These qualities make the StNN an attractive choice for use in resource-constrained contexts where efficiency takes precedence as opposed to computationally expensive data-driven techniques.
V Conclusions
@@
In this paper, we proposed a low-complexity structured neural network (StNN) for modeling and predicting the evolution of dynamical systems starting from the structured matrix theory. Our approach used the Hankel operator to give a structured and computationally efficient alternative to conventional neural networks and data-driven techniques like LEADs, DMD, SINDy, and HAVOK. According to numerical simulations based on the Lotka-Volterra model and the Lorenz system, the proposed StNN outperformed other methods in terms of decreasing complexity, training time, and inference time while retaining accurate long-term trajectory predictions. Our findings show that the structured nature of the Hankel operator-based neural network considerably decreases the number of parameters and flop counts while increasing the efficiency of the StNN when compared to conventional neural networks. Furthermore, comparisons to baseline approaches, such as FFNN, LEADs, DMD, SINDy, and HAVOK demonstrate StNNs’ promise for solving highly nonlinear and chaotic systems with highly accurate long-term predictions.
@@
Future work will include expanding the StNN framework to higher-dimensional dynamical systems and utilizing the Hankel operator defined through the observation of the state to efficiently solve PDEs.